# Linking high-$z$ and low-$z$: Are We Observing the Progenitors of the Milky Way with JWST?

### v1

@@ 2 Methods
Figure 1: Left: Synthetic SED evolution for a MW major branch simulated by the NEFERTITI model, color-coded for redshift. For each SED, we show the corresponding photometries in the JWST NIRCam filters. Right: Star-formation (top panel) and metal-enrichment history (bottom panel) of the same MW major branch. The vertical lines highlight the ages for which the SEDs on the left are shown, with the same color coding. On the bottom-right, we report the $M_{\star}$ at these redshifts.
@@ 2.1 The Milky Way evolution model: NEFERTITI
We employ NEFERTITI (NEar FiEld cosmology: Re-Tracing Invisible TImes, Koutsouridou et al. 2023), a state-of-the-art semi-analytical model for the formation and evolution of a MW-like galaxy that can be coupled with N-body simulations of dark matter (DM) only or with merger trees produced via Monte Carlo algorithms. The model is calibrated to reproduce the present-day observational properties of the MW, like the stellar/gas mass and the metallicity distribution function of halo stars (Bonifacio et al., 2021).
@@ 2.2 Building Up Galaxy Spectra
To make accurate predictions for the emission of the MW progenitors at high-$z$, we build their synthetic SEDs using the spectral synthesis code Yggdrasil (Zackrisson et al., 2011) in the following manner. For each timestep of $1$ Myr, consistently with the NEFERTITI code, we assume an instantaneous star-formation burst. Both the $M_{\star}$ formed and the $Z_{gas}$ from which the stars form, are taken from the output of the galaxy-formation model. With these properties as input, the Yggdrasil code provides the rest-frame spectrum of a simple stellar population (SSP) at different ages (see Zackrisson et al. 2011 for more details). For Pop II stars, the available IMF is the Kroupa 2001 in the interval $m_{\star}=[0.1-100]\,M_{\odot}$ and the SSP are from Starburst99 (Leitherer et al., 1999; Vázquez & Leitherer, 2005). For Pop III stars we choose the lognormal IMF option in the range $m_{\star}=[1-500]\,M_{\odot}$ with the SSP from Raiter et al. 2010. The SEDs shown in this work will have a maximal nebular contribution, thus no escape of Lyman continuum photons, and no dust corrections.
To model the synthetic SED of the progenitor galaxy at a chosen redshift, we assemble the emitted rest-frame spectra of all the star-formation bursts that make up its evolutionary history and shift into the observed frame of reference. Fig. 1 (left) shows the SED evolution of one illustrative MW major branch and the corresponding synthetic photometries in the JWST NIRCam filters. On the right, we show the star-formation rate ($SFR$) and the $Z_{gas}$ as a function of the age of the galaxy (i.e. time since the first star-formation event). The colors mark the ages at which the SED is shown.
@@ 3.1 Detectability of the MW Major Branch
Figure 2: Top: Predictions for the number of detectable progenitors in the virial volume of the MW at different $z$, as a function of the magnification factor. From the 50 merger histories provided by the NEFERTITI model, we calculate the mean number of detections (solid lines) and the standard deviation (shaded areas). Bottom: Comparison between the number of systems detected in the Firefly Sparkle (grey) and the number of progenitors predicted to be detectable in the same conditions (i.e. $z$, volume, lensing). The dashed grey lines highlight the lensing range of the central region of the Firefly Sparkle.
@@ 3.2 Detectability of a MW-type Assembly
Figure 3: Star-formation (gold) and metal-enrichment (blue) histories of the progenitors that best fit the emission of each system of the Firefly Sparkle, in order of increasing estimated $M_{\star}$. The synthetic SED is shown in the top-right inset (red), together with the Mowla et al. 2024 observations (black). The grey areas are not included in fit. The histories of the minor branches (dotted lines) that have merged with the most massive progenitor (solid lines) to form the final one at $z=8.3$ are also shown.
@@ 3.3 Interpreting the Firefly Sparkle
Table 1 reports the physical properties inferred for the 10 Firefly Sparkle systems, which in our interpretation are MW-progenitor galaxies. Their stellar masses are $M_{\star}=10^{6.2-7.5}\,M_{\odot}$, their DM halos have $M_{h}=10^{8.1-9}\,M_{\odot}$, the instantaneous $SFR=0.04-0.20\,M_{\odot}yr^{-1}$, and $Z_{gas}=0.04-0.24\,Z_{\odot}$. We also report $t_{50}$, defined as the time needed to form half of the total stellar mass, which is of the order of $150-300$ Myr. As a result, these galaxies have old stellar populations that have formed more than 100 Myr before the observation.
@@ 4 Discussion and Conclusions
The recent JWST CANUCS observation of the Firefly Sparkle claimed to be a MW-like galaxy assembly at $z\approx 8$, presented us with the unique opportunity to connect the still debated early stages of galaxy formation to our knowledge of the Local Universe. In this Letter, we performed forward modeling of the MW progenitors with the galaxy-formation model NEFERTITI to make both predictions and interpretations for JWST.
First, analyzing multiple realizations of possible MW assembly histories, we find that the most massive MW progenitor can appear within JADES detection limits as early as $z\approx 11$ ($2\%$ probability), it is very likely detectable at $z\approx 8.2$ ($50\%$), and always detectable at $z\leq 7.1$. As a result of this prediction, in both completed and ongoing deep surveys of JWST, we could have observed the major branch of MW-like galaxies that still require interpretation.
@@
Our $SFRs$ are consistent with those inferred by Mowla et al. 2024 with spectrophotometric fitting techniques (Iyer & Gawiser, 2017; Iyer et al., 2019), while their derived $M_{\star}$ are on average lower than our predictions. This is not surprising given the bursty and long SFHs of our MW progenitors, with $\rm t_{50}=150-300$ Myr. Indeed, reconstructing and resolving old stellar populations through SED fitting is particularly difficult due to the outshining of their stellar emission by the dominating nebular emission from young stars (e.g. Whitler et al., 2023; Giménez-Arteaga et al., 2024).
As for the gas metallicity, Mowla et al. (2024) derive a single value for the entire system, $Z_{gas}=0.02\pm 0.01\,Z_{\odot}$. We find the $Z_{gas}$ and its evolution for each MW progenitor, with values that are higher than their estimation but still subsolar. The reason for this discrepancy could be due to the direct $T_{e}$ method used to infer $Z_{gas}$ from the observed spectrum. By fitting the observations with the HOMERUN code (Marconi et al., 2024), which uses a weighted combination of multiple single-cloud photoionization models, we found a total metallicity value that is consistent with those predicted by the NEFERTITI model.
@@ Appendix A Properties of the MW major branches
The 50 merger tree realizations of the NEFERTITI model provide us with 50 different possible formation histories of the MW major branches. Fig. 5 shows the halo mass ($M_{h}$) accretion history and the star-formation history for each merger tree’s most massive MW progenitor, from $z\approx 16$ to $z\approx 6$. The wide range of possible evolutions showcases the importance of using the DM merger tree method (e.g. see Salvadori et al., 2007, for more details), which allows us to have a significant statistical sample. We can notice how the $SFR$ evolves in a stochastic and bursty manner due to the stellar feedback (see Sec. 2.1 and Koutsouridou et al., 2023, for more details on NEFERTITI) and to the galaxy merger process, clearing once more the importance of using physically-motivated SFHs when interpreting JWST observations at high-$z$.
Figure 5: Halo mass accretion history (top panel) and SFH (bottom panel) for all the MW major branches simulated by NEFERTITI, color-coded for the halo mass at $z=6$. The black error bars represent the mean value and the corresponding standard deviation at $z=14,12,10,9,8,7,6$.

### v2

@@ 2 Methods
Figure 1: Left: Average synthetic SEDs for the MW major branch simulated by the NEFERTITI model, color-coded for redshift. For each mean SED, we show the corresponding photometries in the JWST NIRCam filters. The shaded area represents the standard deviation due to the different merger tree realizations. Right: Average star-formation (top panel) and metal-enrichment histories (bottom panel) of the MW major branch, with the same color coding. On the bottom-right, we report the mean $M_{\star}$ at the redshifts shown on the left.
@@ 2.1 The Milky Way evolution model: NEFERTITI
We employ NEFERTITI (NEar FiEld cosmology: Re-Tracing Invisible TImes, Koutsouridou et al. 2023), a state-of-the-art semi-analytical model for the formation and evolution of a MW-like galaxy that can be coupled with N-body simulations of dark matter (DM) only or with merger trees produced via Monte Carlo algorithms.
A MW analog is defined by the following physical properties of our Galaxy at $z=0$: the virial mass $M_{vir}=(1.3\pm 0.3)\times 10^{12}\,M_{\odot}$, the stellar mass $M_{\star}=(5\pm 1)\times 10^{10}\,M_{\odot}$, and the star formation rate $SFR=1-3\,M_{\odot}yr^{-1}$ (e.g. Bland-Hawthorn & Gerhard, 2016), the metallicity distribution function of halo stars (Bonifacio et al., 2021), the gas-to-stellar mass ratio $M_{g}/M_{\star}=0.1-0.15$ (e.g. Ferrière, 2001), and the metallicity of the dense, $Z_{ISM}\approx Z_{\odot}$, and diffuse gas, $Z_{IGM}\approx 0.1-0.3\,Z_{\odot}$ (e.g. Tripp et al., 2003, for more details and references see Sec 2.3 of Koutsouridou et al. 2023.). Indeed, the model is calibrated to reproduce all these present-day observational properties of the MW.
@@ 2.2 Building Up Galaxy Spectra
To make accurate predictions for the emission of the MW progenitors at high-$z$, we build their synthetic SEDs using the spectral synthesis code Yggdrasil (Zackrisson et al., 2011) in the following manner. For each timestep of $1$ Myr, consistently with the NEFERTITI code, we assume an instantaneous instance of star formation. Both the $M_{\star}$ formed and the $Z_{gas}$ from which the stars form, are taken from the output of the galaxy-formation model. With these properties as input, the Yggdrasil code provides the rest-frame spectrum of a simple stellar population (SSP) at different ages (see Zackrisson et al. 2011 for more details). For Pop II stars, the available IMF is the Kroupa 2001 in the interval $m_{\star}=[0.1-100]\,M_{\odot}$ and the SSP are from Starburst99 (Leitherer et al., 1999; Vázquez & Leitherer, 2005). For Pop III stars we choose the lognormal IMF option in the range $m_{\star}=[1-500]\,M_{\odot}$ with the SSP from Raiter et al. 2010. The SEDs shown in this work will have a maximal nebular contribution, thus no escape of Lyman continuum photons, and no dust corrections.
To model the synthetic SED of the progenitor galaxy at a chosen redshift, we assemble the emitted rest-frame spectra of all the star-formation bursts that make up its evolutionary history and shift into the observed frame of reference. Fig. 1 (left) shows the average SED evolution of the MW major branch and the corresponding synthetic photometries in the JWST NIRCam filters. On the right, we show the average $SFR$ and $Z_{gas}$ as a function of the cosmic time. The colors mark the redshifts at which the SED is shown.
@@ 3.1 Detectability of the MW Major Branch
Figure 2: Top: Predictions for the number of detectable progenitors in the virial volume of the MW at different $z$, as a function of the magnification factor. From the 50 merger histories provided by the NEFERTITI model, we calculate the mean number of detections (solid lines) and the standard deviation (shaded areas). Bottom: Comparison between the number of systems detected in the Firefly Sparkle (dark green) and the number of progenitors predicted to be detectable in the same conditions (i.e. $z$, volume, lensing). The white area highlights the lensing range of the central region of the Firefly Sparkle.
@@ 3.2 Detectability of a MW-type Assembly
Figure 3: Star-formation (gold) and metal-enrichment (blue) histories of the progenitors that best fit the emission of each system of the Firefly Sparkle, in order of increasing estimated $M_{\star}$. The synthetic SED is shown in the top-right inset (red), together with the Mowla et al. 2024 observations (black). The grey areas are not included in the fit. The histories of the minor branches (dotted lines) that have merged with the most massive progenitor (solid lines) to form the final one at $z=8.3$ are also shown.
@@ 3.3 Interpreting the Firefly Sparkle
Table 1 reports the physical properties inferred for the 10 Firefly Sparkle systems, which in our interpretation are MW-progenitor galaxies. Their stellar masses are $M_{\star}=10^{6.2-7.5}\,M_{\odot}$, their DM halos have $M_{h}=10^{8.1-9}\,M_{\odot}$, the instantaneous $SFR=0.04-0.20\,M_{\odot}yr^{-1}$, and $Z_{gas}=0.04-0.24\,Z_{\odot}$. We also report $t_{half}$, defined as the time needed to form half of the total stellar mass, which is of the order of $150-300$ Myr. As a result, these galaxies have old stellar populations that have formed more than 100 Myr before the observation.
@@ 4 Discussion and Conclusions
The recent JWST CANUCS observation of the Firefly Sparkle claimed to be a MW-like galaxy assembly at $z\approx 8$, presented us with the unique opportunity to connect the still debated early stages of galaxy formation to our knowledge of the Local Universe. In this Letter, we performed forward modeling of the MW progenitors with the galaxy-formation model NEFERTITI to make both predictions and interpretations for JWST. We chose to use Monte Carlo merger trees to have a valid statistical sample of different possible merger histories for the Milky Way.
First, analyzing the multiple realizations of possible MW assembly histories, we find that the most massive MW progenitor can appear within JADES detection limits as early as $z\approx 11$ ($2\%$ probability), it is very likely detectable at $z\approx 8.2$ ($50\%$), and always detectable at $z\leq 7.1$. As a result of this prediction, in both completed and ongoing deep surveys of JWST, we could have observed the major branch of MW-like galaxies that still require interpretation.
@@
The stellar masses inferred by Mowla et al. 2024 with spectrophotometric fitting techniques (Iyer & Gawiser, 2017; Iyer et al., 2019), $M_{\star,obs}=10^{5.3-5.8}\,M_{\odot}$, are on average lower than our predictions. This is not surprising given the bursty and long SFHs of our MW progenitors, with $\rm t_{half}=150-300$ Myr. Indeed, reconstructing and resolving old stellar populations through SED fitting is particularly difficult due to the outshining of their stellar emission by the dominating nebular emission from young stars (e.g. Whitler et al., 2023; Giménez-Arteaga et al., 2024). However, if we self-consistently calculate the $SFR$ values from the specific $SFRs$ reported in the Firefly Sparkle paper, and their respective errors, we find $SFR_{obs}=0.003-0.6\,M_{\odot}yr^{-1}$, which are consistent with our estimates.
As for the gas metallicity, Mowla et al. (2024) derive a single value for the entire system, $Z_{gas,obs}=0.02\pm 0.01\,Z_{\odot}$. We find the $Z_{gas}$ and its evolution for each MW progenitor, with values that are higher than their estimation but still subsolar. The reason for this discrepancy could be due to the direct $T_{e}$ method used to infer $Z_{gas}$ from the observed spectrum. By fitting the observations with the HOMERUN code (Marconi et al., 2024), which uses a weighted combination of multiple single-cloud photoionization models, we found a total metallicity value that is consistent with those predicted by the NEFERTITI model.
@@ Appendix A Properties of the MW major branches
The 50 merger tree realizations of the NEFERTITI model provide us with 50 different possible formation histories of the MW major branches. Fig. 5 shows the evolution of the physical properties (i.e. $M_{h}$, $M_{\star}$, $SFR$, $Z_{gas}$) for each merger tree’s most massive MW progenitor, from $z\approx 18$ to $z\approx 6$. The wide range of possible evolutions showcases the importance of using the DM merger tree method (e.g. see Salvadori et al., 2007, for more details), which allows us to have a significant statistical sample. We can notice how the $SFR$ evolves in a stochastic and bursty manner due to the stellar feedback (see Sec. 2.1 and Koutsouridou et al., 2023, for more details on NEFERTITI) and to the galaxy merger process, clearing once more the importance of using physically-motivated SFHs when interpreting JWST observations at high-$z$.
Figure 5: Physical properties of all the MW major branches simulated by NEFERTITI: halo mass (top left), stellar mass (bottom left), star formation rate (top right), and gas metallicity (bottom right), colored by halo mass at $z=6$. The black error bars represent the mean value and the relative standard deviation at $z=14,12,10,9,8,7,6$.